### v1

# Flipping operators and locally harmonic Maass forms

Abstract.
In the theory of integral weight harmonic Maass forms of manageable growth, two key differential operators, the Bol operator and the shadow operator, play a fundamental role. Harmonic Maass forms of manageable growth canonically split into two parts, and each operator controls one of these parts. A third operator, called the flipping operator, exchanges the role of these two parts. Maass–Poincaré series (of parabolic type) form a convenient basis of negative weight harmonic Maass forms of manageable growth, and flipping has the effect of negating an index. Recently, there has been much interest in locally harmonic Maass forms defined by the first author, Kane, and Kohnen. These are lifts of Poincaré series of hyperbolic type, and are intimately related to the Shimura and Shintani lifts. In this note, we prove that a similar property holds for the flipping operator applied to these Poincaré series.
Key words and phrases:
Eisenstein series, flipping operator, harmonic Maass forms, integral binary quadratic forms, modular forms, Poincaré series
2020 Mathematics Subject Classification:
11F03, 11F11, 11F25, 11F37
1. Introduction and statement of results
Throughout, we let $k\in\mathbb{N}_{\geq 2}$. The flipping operator is defined as (see [4])
$\mathfrak{F}_{2-2k}(f)(\tau)\coloneqq-\frac{v^{2k-2}}{(2k-2)!}\overline{R_{2-2k}^{2k-2}(f)(\tau)},$
where $\tau=u+iv\in\mathbb{H}$ throughout. Here, for $\kappa\in\mathbb{Z}$, the (iterated) Maass raising operator is given as ($n\in\mathbb{N}$)
$R_{\kappa}\coloneqq 2i\frac{d}{d\tau}+\frac{\kappa}{v},\quad R_{\kappa}^{n}\coloneqq R_{\kappa+2(n-1)}\circ\ldots\circ R_{\kappa}.$
The goal of this paper is to show that the flipping operator keeps a certain quadratic form Poincaré series invariant up to sign.
The flipping operator acts on the weight $2-2k$ Maass–Poincaré series $\mathcal{P}_{2-2k,m}$ of index $m\in\mathbb{Z}\setminus\{0\}$ defined in (2.2) by (see [2, Theorem 6.11 iv)])
$\mathfrak{F}_{2-2k}(\mathcal{P}_{2-2k,m})=\mathcal{P}_{2-2k,-m}.$
This “flipping” of $m$ and $-m$ is actually where the name “flipping operator” comes from. For the Maass–Eisenstein series $\mathcal{E}_{2k}$ defined in (2.3), we have (see [2, Theorem 6.15 iv)])
$\mathfrak{F}_{2-2k}(\mathcal{E}_{2k})=-\mathcal{E}_{2k}.$ (1.1)
The flipping operator is closely related to the Bol operator [6] and the shadow operator [5], which are respectively given by
$\mathcal{D}^{2k-1}\coloneqq\left(\frac{1}{2\pi i}\frac{\partial}{\partial\tau}\right)^{2k-1},\qquad\xi_{2-2k}\coloneqq 2iv^{2-2k}\overline{\frac{\partial}{\partial\overline{\tau}}}.$
Let $P_{2k,m}$ and $E_{2k}$ be the usual (weakly) holomorphic Poincaré and Eisenstein series, respectively, defined in (2.1). Then (see [2, Theorem 6.11 ii), iii))], we have
$\xi_{2-2k}(\mathcal{P}_{2-2k,m})=-\frac{(4\pi m)^{2k-1}}{(2k-2)!}P_{2k,-m},\qquad\mathcal{D}^{2k-1}(\mathcal{P}_{2-2k,m})=m^{2k-1}P_{2k,m}.$
For the Maass–Eisenstein series $\mathcal{E}_{2-2k}$ (see [2, Theorem 6.15 ii), iii)]), we have
$\xi_{2-2k}(\mathcal{E}_{2-2k})=(2k-1)E_{2k},\quad\mathcal{D}^{2k-1}(\mathcal{E}_{2-2k})=-(4\pi)^{1-2k}(2k-1)!E_{2k}.$ (1.2)
We refer to (2.7) and (2.8) for the interplay between $\mathfrak{F}_{2-2k}$, $\xi_{2-2k}$, and $\mathcal{D}^{2k-1}$.
The aim of this paper is to show that (1.1) extends to a certain quadratic form Poincaré series. To describe this, we let $D>0$ be a non-square discriminant, and $\mathcal{Q}_{D}$ be the set of integral binary quadratic forms $Q(x,y)\coloneqq[a,b,c](x,y)\coloneqq ax^{2}+bxy+cy^{2}$ of discriminant $D\coloneqq b^{2}-4ac$ throughout. Zagier [17] defined
$f_{k,D}(\tau)\coloneqq\frac{D^{k-\frac{1}{2}}}{\binom{2k-2}{k-1}\pi}\sum_{Q\in\mathcal{Q}_{D}}\frac{1}{Q(\tau,1)^{k}}\in S_{2k},$
where $S_{2k}$ denotes the space of weight $2k$ cusp forms for $\mathrm{SL}_{2}(\mathbb{Z})$. These functions have very fruitful applications in the theory of modular forms [12, 14, 15]. Katok [13] showed that one can write $f_{k,D}$ as a hyperbolic Eisenstein series (see [11] for more details on hyperbolic expansions), and used this to prove that the $f_{k,D}$ generate $S_{2k}$ as $D$ ranges over positive discriminants. Paralleling the behavior of the parabolic Eisenstein series in (1.2), the first author, Kane, and Kohnen [3] discovered a weight $2-2k$ preimage $\mathcal{F}_{1-k,D}$ of $f_{k,D}$ (up to constants) under both the Bol operator and the shadow operator . To introduce it, we define for $Q=[a,b,c]\in\mathcal{Q}_{D}$
[…]
for $x\in[0,1]$ and $r,s>0$. Then the function $\mathcal{F}_{1-k,D}$ is given by (see [3, (1.4)])
$\mathcal{F}_{1-k,D}(\tau)\!\coloneqq\!\frac{D^{\frac{1}{2}-k}}{2\binom{2k-2}{k-1}\pi}\sum_{Q\in\mathcal{Q}_{D}}\operatorname{sgn}(Q_{\tau})Q(\tau,1)^{k-1}\beta\!\left(\frac{Dv^{2}}{\left\lvert Q(\tau,1)\right\rvert^{2}};k\!-\!\frac{1}{2},\frac{1}{2}\right)\!.$ (1.3)
Analogous to $f_{k,D}$, one may view $\mathcal{F}_{1-k,D}$ as a hyperbolic Eisenstein series of negative weight. Paralleling (1.2), it was shown in [3, Theorem 1.2],
$\xi_{2-2k}(\mathcal{F}_{1-k,D})=D^{\frac{1}{2}-k}f_{k,D},\quad\mathcal{D}^{2k-1}(\mathcal{F}_{1-k,D})=-\frac{(2k-2)!}{(4\pi)^{2k-1}}D^{\frac{1}{2}-k}f_{k,D}$ (1.4)
outside a certain “exceptional set” $E_{D}\subseteq\mathbb{H}$ (see (2.9)) of measure [^1] The set $E_{D}$ is a union of geodesics associated to $Q\in\mathcal{Q}_{D}$. 00. Thus, it is natural to ask if (1.1) extends from a parabolic to a hyperbolic setting as well. We show that this is indeed the case.
Theorem 1.1.
If $D\in\mathbb{N}$ is a non-square discriminant, $k\geq 2$, and $\tau\not\in E_{D}$, then we have
$\mathfrak{F}_{2-2k}(\mathcal{F}_{1-k,D})(\tau)=-\mathcal{F}_{1-k,D}(\tau).$
Remark.
In [9, 16], the function $\mathcal{F}_{1-k,D}$ played a crucial role to characterize non-trivial vanishsing of twisted central $L$-values. Theorem 1.1 might help to simplify the methods developed there.
The paper is organized as follows. In Section 2 we review preliminaries on modular forms, harmonic Maass forms, and locally harmonic Maass forms as well as their properties. We conclude by proving Theorem 1.1 in Section 3.
Acknowledgments
The first author was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001179). The second author received support for this research by an AMS–Simons Travel Grant. The third author’s work was supported by a grant from the Simons Foundation (853830, LR).
2. Preliminaries
We summarize some standard notions from the theory of holomorphic modular forms and harmonic Maass forms. We refer to [2] for more details.
2.1. Poincaré series
For a function $f\colon\mathbb{H}\rightarrow\mathbb{C}$, $\kappa\in\mathbb{Z}$, and a matrix $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma\coloneqq\mathrm{SL}_{2}(\mathbb{Z})$, define the Petersson slash operator
$f\big{|}_{\kappa}\begin{pmatrix}a&b\\ c&d\end{pmatrix}(\tau)\coloneqq(c\tau+d)^{-\kappa}f\left(\frac{a\tau+b}{c\tau+d}\right).$
For completeness, we define the variants of modular forms we need.
Definition 2.1.
Let $f\colon\mathbb{H}\to\mathbb{C}$ be a function and $\kappa\in\mathbb{Z}$.
  1. (1)
    We say $f$ is a holomorphic modular form of weight $\kappa$ if the following hold:
    1. (i)
      For $\gamma\in\Gamma$ we have $f|_{\kappa}\gamma=f$,
    2. (ii)
      $f$ is holomorphic on $\mathbb{H}$,
    3. (iii)
      $f$ is holomorphic at $i\infty$.
    We denote the vector space of functions satisfying these conditions by $M_{\kappa}$.
  2. (2)
    If in addition $f$ vanishes at $i\infty$, then we call $f$ a cusp form. The space of cusp forms is denoted by $S_{\kappa}$.
  3. (3)
    If $f$ satisfies the conditions (i) and (ii) from (1) and is permitted to have a pole at $i\infty$, then we call $f$ a weakly holomorphic modular form of weight ${\kappa}$. The vector space of such functions is denoted by $M_{\kappa}^{!}$.
To describe canonical examples, for $k\in\mathbb{N}_{\geq 2}$ and $m\in\mathbb{Z}$, set
$P_{2k,m}\coloneqq\sum_{\gamma\in\Gamma_{\infty}\backslash\Gamma}\varphi_{m}\big{|}_{2k}\gamma,\qquad E_{2k}\coloneqq P_{2k,0},$ (2.1)
where $\varphi_{m}(\tau)\coloneqq q^{m}$ with $q\coloneqq e^{2\pi i\tau}$ and $\Gamma_{\infty}\coloneqq\left\{\pm\begin{pmatrix}1&n\\ 0&1\end{pmatrix}\colon n\in\mathbb{Z}\right\}$. We have
$P_{2k,m}\in\begin{cases}S_{2k}&\text{if }m>0,\\ M_{2k}&\text{if }m=0,\\ M_{2k}^{!}&\text{if }m<0,\end{cases}$
and $P_{2k,m}$ spans $S_{2k}$ resp. $M_{2k}^{!}$ for $m\neq 0$ (see [2, Theorems 6.8 and 6.9]).
For negative $k$, we require the following non-holomorphic modular forms.
Definition 2.2.
Let $f\colon\mathbb{H}\to\mathbb{C}$ be a smooth function.
  1. (1)
    We say $f$ is a weight $2-2k$ harmonic Maass form if the following hold:
    1. (i)
      For $\gamma\in\Gamma$ we have $f|_{2-2k}\gamma=f$.
    2. (ii)
      We have $\Delta_{2-2k}(f)=0,$ where the weight $\kappa$ hyperbolic Laplacian is
      $\Delta_{2-2k}\coloneqq-v^{2}\left(\frac{\partial^{2}}{\partial u^{2}}+\frac{\partial^{2}}{\partial v^{2}}\right)+i(2-2k)v\left(\frac{\partial}{\partial u}+i\frac{\partial}{\partial v}\right).$
    3. (iii)
      There exists a polynomial $P_{f}(\tau)\in\mathbb{C}[q^{-1}]$, the principal part of $f$ such that
      $f(\tau)-P_{f}(\tau)=O\left(e^{-\varepsilon v}\right)$
      as $v\to\infty$ for some $\varepsilon>0$.
    We denote the vector space of such functions by $H_{\kappa}$.
  2. (2)
    If $f$ satisfies the conditions (i) and (ii) from (1) as well as
    $f(\tau)=O\left(e^{\varepsilon v}\right)\ \text{as}\ v\to\infty\ \text{for some}\ \varepsilon>0,$
    then we call $f$ a harmonic Maass form of manageable growth. We denote the vector space of such forms by $H_{\kappa}^{!}$.
Let $M_{\mu,\nu}$ be the $M$-Whittaker function (see [1, Subsection 13.14] for example) and define the Maass–Poincaré series (see [2, Definition 6.10])
$\mathcal{P}_{2-2k,m}\coloneqq\sum_{\gamma\in\Gamma_{\infty}\backslash\Gamma}\phi_{2-2k,m}\big{|}_{2-2k}\gamma,\qquad m\in\mathbb{Z}\setminus\{0\},$ (2.2)
where
$\phi_{2-2k,m}(\tau)\coloneqq\frac{(-\operatorname{sgn}(m))^{2k-1}(4\pi|m|v)^{k-1}}{(2k-1)!}M_{\operatorname{sgn}(m)(1-k),k-\frac{1}{2}}(4\pi|m|v)e^{2\pi imu}.$
The Maass–Eisenstein series is given by (see [2, p. 104])
$\mathcal{E}_{2-2k}\coloneqq\sum_{\gamma\in\Gamma_{\infty}\backslash\Gamma}v^{2k-1}\big{|}_{2-2k}\gamma.$ (2.3)
We have $\mathcal{E}_{2-2k}\in H_{2-2k}^{!}$ and
$\mathcal{P}_{2-2k,m}\in\begin{cases}H_{2-2k}^{!}&\text{if }m>0,\\ H_{2-2k}&\text{if }m<0.\end{cases}$
By virtue of their Fourier expansion (see [2, Theorem 6.11 v)]), $\mathcal{P}_{2-2k,m}$ has a prescribed prinicpal part for $m<0$ resp. a prescribed non-holomorphic part for $m>0$ (see (2.5) resp. (2.6) below). For this reason, $\mathcal{P}_{2-2k,m}$ span $H_{2-2k}$ (resp. $H_{2-2k}^{!}$ if restricting to $m>0$). We refer the reader to [2, Theorems 6.11 and 6.15] for proofs including (1.1), (1.2) and more details.
2.2. Differential operators and Fourier expansions
Let $\gamma\in\Gamma$ and $f\colon\mathbb{H}\to\mathbb{C}$ be smooth. Then, the Maass raising operator satisfies (see [2, Lemma 5.2])
$R_{\kappa}\left(f|_{\kappa}\gamma\right)=R_{\kappa}(f)|_{\kappa+2}\gamma.$
Moreover, if $f$ is an eigenfunction of $\Delta_{\kappa}$ with eigenvalue $\lambda$, then $R_{\kappa}(f)$ is an eigenfunction of $\Delta_{\kappa+2}$ with eigenvalue $\lambda+\kappa$, see [2, Lemma 5.2] as well. Furthermore, according to [7, (56)], the iterated Maass raising operator and the Bol operator are related by
$R_{2-2k}^{n}=\sum_{r=0}^{n}(-1)^{r}\binom{n}{r}(2-2k+r)_{n-r}v^{r-n}(4\pi)^{r}\mathcal{D}^{r},$ (2.4)
where the rising factorial is defined as
$(a)_{n}\coloneqq a(a+1)\cdots(a+n-1),\qquad n\in\mathbb{N}_{0}.$
Bruinier and Funke [5] showed that the Fourier expansion of a harmonic Maass form of manageable growth [^2] The reader should be aware that their terminology refers to our harmonic Maass forms of manageable growth as “weak Maass forms.” $f\in H_{2-2k}^{!}$ naturally splits into a holomorphic part and a non-holomorphic part. Namely (see [2, Lemma 4.3]), we have a Fourier expansion of the shape
$f(\tau)=\sum_{n\gg-\infty}c_{f}^{+}(n)q^{n}+c_{f}^{-}(0)v^{2k-1}+\sum_{\begin{subarray}{c}n\ll\infty\\ n\neq 0\end{subarray}}c_{f}^{-}(n)\Gamma(2k-1,-4\pi nv)q^{n}.$ (2.5)
Here, the non-holomorphic part involves the incomplete Gamma function
[…]
defined for $\operatorname{Re}(s)>0$ and $x\in\mathbb{R}$. In particular, if $f\in H_{2-2k}$ then $f$ has a Fourier expansion of the shape
$f(\tau)=\sum_{n\gg-\infty}c_{f}^{+}(n)q^{n}+\sum_{n<0}c_{f}^{-}(n)\Gamma(2k-1,-4\pi nv)q^{n}.$ (2.6)
By [2, Proposition 5.15 iii), iv)], the flipping operator satisfies
$\xi_{2-2k}(\mathfrak{F}_{2-2k}(f))$ $=\frac{(4\pi)^{2k-1}}{(2k-2)!}\mathcal{D}^{2k-1}(f),$ (2.7)
$\mathcal{D}^{2k-1}(\mathfrak{F}_{2-2k}(f))$ $=\frac{(2k-2)!}{(4\pi)^{2k-1}}\xi_{2-2k}(f).$ (2.8)
In other words, the flipping operator “switches” the holomorphic and non-holomorphic part in the Fourier expansion of a harmonic Maass form of manageable growth.
2.3. Locally harmonic Maass forms
Let
$E_{D}\coloneqq\bigcup_{Q\in\mathcal{Q}_{D}}\left\{\tau\in\mathbb{H}\colon Q_{\tau}=0\right\}.$ (2.9)
We recall the following definition from [3, Section 2].
Definition 2.3.
A function $f\colon\mathbb{H}\to\mathbb{C}$ is called a locally harmonic Maass form of weight $2-2k$ with exceptional set $E_{D}$, if it satisfies the following conditions:
  1. (1)
    We have $f|_{2-2k}\gamma=f$ for every $\gamma\in\Gamma$.
  2. (2)
    For all $\tau\in\mathbb{H}\setminus E_{D}$, there exists a neighborhood of $\tau$, in which $f$ is real-analytic and $\Delta_{2-2k}(f)(\tau)=0$.
  3. (3)
    For every $\tau\in E_{D}$, we have that
    $f(\tau)=\frac{1}{2}\lim_{\varepsilon\to 0^{+}}(f(\tau+i\varepsilon)+f(\tau-i\varepsilon)).$
  4. (4)
    The function $f$ exhibits at most polynomial growth towards $i\infty$.
According to [3, Theorem 1.1], $\mathcal{F}_{1-k,D}$ from (1.3) is such a locally harmonic Maass form with exceptional set $E_{D}$. A key property of $\mathcal{F}_{1-k,D}$ is that it admits a certain splitting, which might be viewed as a generalization of the Fourier expansion (2.5). To describe this, we define the holomorphic and non-holomorphic Eichler integrals [10, Section 2] of a cusp form $f(\tau)=\sum_{n\geq 1}c_{f}(n)q^{n}\in S_{2k}$ by
[…]
where $f^{c}(z)=\overline{f(-\overline{z})}$. We have
$\xi_{2-2k}\left(f^{*}\right)=f,\quad\mathcal{D}^{2k-1}\left(f^{*}\right)=0,\quad\xi_{2-2k}\left(\mathcal{E}_{f}\right)=0,\quad\mathcal{D}^{2k-1}\left(\mathcal{E}_{f}\right)=f.$
Note that both $\mathcal{E}_{f}$ and $f^{*}$ are real-analytic for all $\tau\in\mathbb{H}$.
Lemma 2.4 ([3, Theorem 7.1]).
Let $\mathcal{C}\subseteq\mathbb{H}\setminus E_{D}$ be a connected component and $\tau\in\mathcal{C}$. Let
$c_{\infty}\coloneqq-\frac{1}{2^{2k-2}(2k-1)\binom{2k-2}{k-1}}\sum_{a\geq 1}\sum_{\begin{subarray}{c}0\leq b<2a\\ b^{2}\equiv D\pmod{4a}\end{subarray}}\frac{1}{a^{k}},$
and define the local polynomial by
$P_{\mathcal{C}}(\tau)\coloneqq c_{\infty}+(-1)^{k}2^{3-2k}D^{\frac{1}{2}-k}\sum_{\begin{subarray}{c}Q=[a,b,c]\in\mathcal{Q}_{D}\\ a<0<Q_{\tau}\end{subarray}}Q(\tau,1)^{k-1}.$ (2.10)
Then, we have
$\mathcal{F}_{1-k,D}(\tau)=D^{\frac{1}{2}-k}f_{k,D}^{*}(\tau)-D^{\frac{1}{2}-k}\frac{(2k-2)!}{(4\pi)^{2k-1}}\mathcal{E}_{f_{k,D}}(\tau)+P_{\mathcal{C}}(\tau).$
Remark.
The constant $c_{\infty}$ can be evaluated using [18, Proposition 3].
Note that the splitting in Lemma 2.4 resembles the Fourier expansion (2.5) of a harmonic Maass form of manageable growth with $c_{f}^{-}(0)v^{k-1}$ replaced by the local polynomial $P_{\mathcal{C}}$.
3. Proof of Theorem 1.1
In this section we prove our main result.
Proof of Theorem 1.1.
We use Lemma 2.4 and define the real-analytic function
$\mathbb{F}\coloneqq D^{\frac{1}{2}-k}f_{k,D}^{*}-D^{\frac{1}{2}-k}\frac{(2k-2)!}{(4\pi)^{2k-1}}\mathcal{E}_{f_{k,D}}.$
Using (2.7), (2.8), and (1.4) gives that
$\xi_{2-2k}(\mathfrak{F}_{2-2k}(\mathbb{F}))=-\xi_{2-2k}(\mathbb{F}),\quad\mathcal{D}^{2k-1}(\mathfrak{F}_{2-2k}(\mathbb{F}))=-\mathcal{D}^{2k-1}(\mathbb{F}).$
Since the first term of $\mathbb{F}$ decays towards $i\infty$ and its second term is a $q$-series without a constant term, this shows that
$\mathfrak{F}_{2-2k}(\mathbb{F})=-\mathbb{F}.$
So we are left to show that $\mathfrak{F}_{2-2k}(P_{\mathcal{C}})=-P_{\mathcal{C}}$. By definition of $P_{\mathcal{C}}$ in (2.10), it suffices to show that
$\mathfrak{F}_{2-2k}(1)$ $=-1,$ (3.1)
$\mathfrak{F}_{2-2k}(p)$ $=-p,$ (3.2)
where
$p(\tau)\coloneqq(-1)^{k}2^{3-2k}D^{\frac{1}{2}-k}\sum_{\begin{subarray}{c}Q=[a,b,c]\in\mathcal{Q}_{D}\\ a<0<Q_{\tau}\end{subarray}}Q(\tau,1)^{k-1}.$
We start by proving (3.1), which is equivalent to showing that
$R_{2-2k}^{2k-2}(1)=(2k-2)!v^{2-2k}.$
This follows directly using (2.4).
We next prove (3.2). Since $\tau\not\in E_{D}$ and $D$ a non-square, we may rewrite
$p(\tau)=(-1)^{k}2^{1-2k}D^{\frac{1}{2}-k}\sum_{Q=[a,b,c]\in\mathcal{Q}_{D}}\left(1-\operatorname{sgn}(a)\right)\left(\operatorname{sgn}\left(Q_{\tau}\right)+1\right)Q(\tau,1)^{k-1}.$
Next observe that
$R_{\kappa}\left(\left(\operatorname{sgn}\left(Q_{\tau}\right)+1\right)Q(\tau,1)^{k-1}\right)=\left(\operatorname{sgn}\left(Q_{\tau}\right)+1\right)R_{\kappa}\left(Q(\tau,1)^{k-1}\right)$
for $\kappa\in\mathbb{Z}$, because $\frac{d}{d\tau}(\operatorname{sgn}(Q_{\tau})+1)=0$. Thus, (3.2) follows if we show
$\mathfrak{F}_{2-2k}\!\left(Q(\tau,1)^{k-1}\right)=-Q(\tau,1)^{k-1}.$
This is equivalent to
$R_{2-2k}^{2k-2}\!\left(Q(\tau,1)^{k-1}\right)=\frac{(2k-2)!}{v^{2k-2}}Q(\overline{\tau},1)^{k-1}.$ (3.3)
To prove this note that there exists $A\in\mathrm{SL}_{2}(\mathbb{R})$ (see [3, Lemma 3.1]) such that
$\tau|_{-2}A=-\frac{Q(\tau,1)}{\sqrt{D}}.$ (3.4)
By the definition of the slash operator, (3.4) is equivalent to
$\tau^{k-1}\big{|}_{2-2k}A=\frac{(-1)^{k+1}}{D^{\frac{k-1}{2}}}Q(\tau,1)^{k-1}.$
By [8, Lemma 2.1.1], the raising operator and the slash operator commute on $\mathrm{SL}_{2}(\mathbb{R})$. Thus, we have
$R_{2-2k}^{2k-2}\left(Q(\tau,1)^{k-1}\right)=(-1)^{k+1}D^{\frac{k-1}{2}}R_{2-2k}^{2k-2}\left(\tau^{k-1}\right)\Big{|}_{2k-2}A.$
To prove (3.3), and hence (3.2) as well, we next claim that it suffices to verify
$R_{2-2k}^{2k-2}\left(\tau^{k-1}\right)=(2k-2)!\left(\frac{\overline{\tau}}{v^{2}}\right)^{k-1}.$ (3.5)
Indeed, assuming (3.5) gives
$R_{2-2k}^{2k-2}\left(Q(\tau,1)^{k-1}\right)=(-1)^{k+1}D^{\frac{k-1}{2}}(2k-2)!\left(\frac{\overline{\tau}^{k-1}}{v^{2k-2}}\right)\bigg{|}_{2k-2}A.$
Rewriting yields
$\left(\frac{\overline{\tau}^{k-1}}{v^{2k-2}}\right)\bigg{|}_{2k-2}A=\frac{(-1)^{k+1}}{D^{\frac{k-1}{2}}}v^{2-2k}Q(\overline{\tau},1)^{k-1}.$
This directly gives (3.3).
Hence, we are left to prove (3.5). By (2.4), we have
$R_{2-2k}^{2k-2}\left(\tau^{k-1}\right)=\sum_{r=0}^{2k-2}(-1)^{r}\binom{2k-2}{r}(2-2k+r)_{2k-2-r}v^{r-2k+2}(4\pi)^{r}\mathcal{D}^{r}\left(\tau^{k-1}\right).$
If $r>k-1$, then $\mathcal{D}^{r}(\tau^{k-1})=0$. If $r\leq k-1$, then
$\mathcal{D}^{r}\left(\tau^{k-1}\right)=\frac{1}{(2\pi i)^{r}}\frac{(k-1)!}{(k-r-1)!}\tau^{k-1-r}.$
Moreover
$(2-2k+r)_{2k-2-r}=(-1)^{r}(2k-r-2)!.$
We then verify (3.5) by obtaining
$R_{2-2k}^{2k-2}\!\left(\tau^{k-1}\right)$ $=(2k\!-\!2)!\!\left(\frac{\tau}{v^{2}}\right)^{\!k-1}\sum_{r=0}^{k-1}\frac{(k\!-\!1)!}{r!(k\!-\!1\!-\!r)!}\!\left(\frac{-2iv}{\tau}\right)^{\!r}=(2k\!-\!2)!\!\left(\frac{\overline{\tau}}{v^{2}}\right)^{\!k-1}.$
This completes the proof. ∎